OpenStudy (anonymous):

Can someone check my work for me? Given Triangle ABC, find the length of AB. Which theorem did you use?

OpenStudy (anonymous):



OpenStudy (anonymous):

i used the Pythagorean theorem and got 12

OpenStudy (jacobchaotic):

Do you happen to have a list of available answers?

OpenStudy (anonymous):

@JacobChaotic It's not multiple choice, I have to solve it

OpenStudy (anonymous):

you use pythagorean theorem

OpenStudy (anonymous):

15^2-9^2 is the equation you need to use

OpenStudy (anonymous):

then you will need to take the square root of the answer

OpenStudy (anonymous):

12 is the correct answer

OpenStudy (anonymous):

I thought the Pythagorean theorem was a2+b2=c2 ? I did 9^2+b^2=15^2 and still ended up with 12

OpenStudy (jacobchaotic):

Oh man, that's terrible. I sort of have trouble with the Pythagorean Theorem, but I believe the answer would be 12 also. The hypotenuse (15) of the triangle would be C. Then, you could either apply 9 for A or B. For this problem, I'm gonna replace A with 9. Pythagorean Theorem: a^2+b^2=c^2 9^2+6^2=15^2 81+36=225 117=225 Then, you would divide 225 by 117 getting you 1.92 which means they're almost half of each other. So, 6 times 1.92 = 11.52, which you could round the answer to 12.

OpenStudy (anonymous):

a^2+b^2=c^2 is for working out the hypotenuse. for other sides you use c^2-b^2=a^2

OpenStudy (anonymous):

it isn't 6^2. it is 12^2

OpenStudy (anonymous):

so i do 15^2-9^2 and then 225-81=144 and take the square root which is 12

OpenStudy (loser66):

you are correct

OpenStudy (anonymous):

ok thank you :)